OpenStudy (anonymous):

Solve by elimination 9x^2 + y^2 = 85 2x^2 - 3y^2 = 6

OpenStudy (uri):

Mutliply 1st eq with 3

OpenStudy (uri):

What do you get?

OpenStudy (anonymous):

So 1st equ. is 27x^3+ 3y^2 = 225

OpenStudy (anonymous):

2x^2 + 27x^3 = 231

OpenStudy (uri):

27x^3 o_O?

OpenStudy (anonymous):

opps 29x^2

OpenStudy (anonymous):

the same power

OpenStudy (uri):

Now divide 231 with 29

OpenStudy (anonymous):

7.96 take the square root 2.822

OpenStudy (uri):

Wait you mutilpied wrong...

OpenStudy (uri):

9x^2 + y^2 = 85 2x^2 - 3y^2 = 6 First eq mutiply by 3 will be 27x^2+3y^2=255

OpenStudy (uri):

So, 27x^2+3y^2=255 2x^2 - 3y^2 = 6 =========== 29x^2=261 divide 261 with 29 x^2=9

OpenStudy (anonymous):

i see ...9 take the square root x=3

OpenStudy (uri):

Ya!

OpenStudy (anonymous):

substitute the x value to find y right?

OpenStudy (uri):

Yes! in the ORIGINAL second eq i.e 2x^2 - 3y^2 = 6

OpenStudy (anonymous):

y = square root of 10

OpenStudy (uri):

2(9)-3y^2=6 18-3y^2=6 subract 18 both sides 18-3y^2=6 -18 -18 ========= -3y^2=-12 y^2=-12/-3 y^2=4

OpenStudy (uri):

y=2